OpenStudy (anonymous):

Literal Equation: Help solve 3x-4y = 26 x - 8y = 22

OpenStudy (amistre64):

3x-4y = 26 (-2) x - 8y = 22 -6x+8y = -52 x - 8y = 22 ------------- -5x = -30 x = 6

OpenStudy (anonymous):

x=6,y=-2

OpenStudy (anonymous):

Why do you only subtract -2?

OpenStudy (anonymous):

he is multiplying the whole equation with -2

OpenStudy (amistre64):

you mulltiply the whole thing by (-2) in order to get an equation that can eliminate a variable

OpenStudy (anonymous):

And you are elimination the Y?

OpenStudy (amistre64):

yes; with the y gone, we can easily solve for x

OpenStudy (anonymous):

What about substitution?

OpenStudy (amistre64):

for example: 3x = 12; x=4 what happens if we multiply the whole thing by 5, does the x value change? 3x = 12 (5) 15x = 60 ; x = 4

OpenStudy (amistre64):

substitution is just solving for one variable in the terms of the other ... its more labor intensive

OpenStudy (amistre64):

subbing, elimiation, matrix .... all get the same answer so you eventually just go with whats best for you :)

OpenStudy (anonymous):

Except rumor has it that this problem will be on a test with subbing instead!

OpenStudy (amistre64):

How do you get to kentucky? lol

OpenStudy (amistre64):

then solve for one variable in terms of the other; use that 'value' in the other eqaution to get a solid value for the 'other' variable; and use the solid value to determine the value of the unknowned

OpenStudy (anonymous):

What if i took 3x - 4y = 26 x - 8y = 22 and i did this x = 22 + 8y would that work?

OpenStudy (amistre64):

that would work out great :)

OpenStudy (amistre64):

since x=x=x=x=x... use that 'value' for x in the top and solve it all for y

OpenStudy (anonymous):

Are you saying 3(22 + 8y) - 4y = 26 66 + 24y - 4y = 26

OpenStudy (amistre64):

thats correct

OpenStudy (anonymous):

Why do I flip it the 2nd equation though?

OpenStudy (amistre64):

becasue you cant really determine the value of y from the first equation; you need another 'form' to use it in to calibrate the value

OpenStudy (amistre64):

umm.... what do you mean by flip it?

OpenStudy (anonymous):

Changing it from x - 8y = 22 to x = 22 + 8y

OpenStudy (amistre64):

think of subbing as 'looking at it from another direction'...

OpenStudy (amistre64):

x + y = 6 can have any values; and you cant determine x or y to be one point ..... so look at it thru a different equation right?

OpenStudy (anonymous):

So in the x - y = 4 4y - x = 14 could I change the x - y = 4 to x = 4 + y?

OpenStudy (amistre64):

yes..... then we use that 'value' in the other equation to see it from a different perspective and can solve for y in it

OpenStudy (anonymous):

or in the equation x + y = 10 6x - 7y = 34 I can change the x + y = 10 to x = 10 - y?

OpenStudy (amistre64):

exactly

OpenStudy (anonymous):

I think I get it. If I don't I'll just ask you again :D

OpenStudy (amistre64):

good idea ;)